OpenStudy (anonymous):

Will give medal and become fan!!! Identify the 35th term of an arithmetic sequence where a1 = -7 and a18 = 95. Answer Choices 197 163 203 168

OpenStudy (anonymous):

@arjund12

OpenStudy (anonymous):

@fbi2015

OpenStudy (anonymous):

@brooke2353

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@ayanna1100

OpenStudy (bohotness):

yes ?

OpenStudy (brooke2353):

a18=a1+(n-1)*d 95=-7+17d 17d=102 d=6 a35=a1+(n-1)*d=-7+34*6=197

OpenStudy (brooke2353):

I hope that helps

OpenStudy (anonymous):

@brooke2353 thanks so much!

OpenStudy (brooke2353):

mhm its no problem that was really easy for me